The University of Texas at Austin

ITS Services Status

Current Service Interruptions

5/5/2016
3:18 PM
RESOLVED - SQL 2012 unplanned maintenance
Service Restored

05/05/2016 3:09
SQL2012 unplanned maintenance has concluded.


05/05/2016 2:43 PM
One of the SQL 2012 Production instances is having performance issues.  ITS Systems Database team is currently working on resolving the issue. 
Customers who are connecting to SAGL1, SAGL2 and SAGL3 will experience a brief connection interruptions while we migrate the Availability Groups to other SQL instances.

5/4/2016
3:41 PM
Intermittent Outage for some UCDs (Universal Call Distribution)
Service Functioning, Temporary Solution in Place

5/4/2016 3:39 PM
The workaround has been applied to all know UCDs. If other UCD customers experience this issue, please contact assignment@its.utexas.edu as soon as possible.


5/4/2016 11:50 AM
Some UCD (Universal Call Distribution) customers are reporting an issue with UCDs going to overflow routes and not to agents who are logged in.  ITS is implementing a workaround for those affected UCD groups and will be engaging the vendor. 

4/15/2016
11:45 PM
Hitachi NAS Configuration Issue
Unplanned Maintenance

4/15/2016 11:41 PM
Mitigation efforts are ongoing. The majority of clients have been updated and the rest systems are waiting on maintenance windows.


3/19/2016 9:00 AM
The mitigation effort is underway. Next update will be on 3/25


3/12/16 9pm
The Enterprise Systems Management team is devising the plan for NFS connection string mitigation.  Next update Friday Mar 18, 2016.


3/4/16 4:48pm
Current Status: Enterprise Service Management team continues to plan the NFS connect string mitigation effort.

Next Update: EOD March 11, 2016


3/2/2016 4:56pm
What: The Hitachi NAS (HNAS) storage array is running in an unanticipated configuration following a recent maintenance event.

Impact: There is no known impact. There is no anticipated impact from mitigation efforts that are initiating now.

Scope of Impact: All devices/services that mount storage from the HNAS are in scope.

Mitigation: The Enterprise Systems Management team is working with the Storage team to update NFS connect strings on all servers in scope.

Next Update: EOD Friday March 4, 2016.

Past Service Interruptions

Upcoming Maintenance

<<< 02/26/2012 - 03/17/2012 >>>

  Today's date

We Can Help

Get help from an expert:

* ITS Help and Service Desk

* Call us at 512-475-9400

* Submit a help request online

We also have a walk-in service in the first floor lobby of the Flawn Academic Center (FAC). Stop by and let us help you!